OpenStudy (anonymous):

N(t)=0.81t-1.14square root t + 1.53. Find absolute extrema on [0,6. derivative -0.19 - 0.57/t^.5 the only critical value is 0 (right? check please) I found N(6) = 3.60 The answers page also has N(.5) I don't see where the came from.

OpenStudy (anonymous):

\[N(t)=.81t-1.14\sqrt{t+1.53}\] i hate decimals

OpenStudy (anonymous):

or did i get the function wrong?

OpenStudy (anonymous):

the square root is only over the t

OpenStudy (anonymous):

\[N(t)=.81t-1.14\sqrt{t}+1.53\]

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

derivative is \[.81-\frac{.57}{\sqrt{t}}\]

OpenStudy (anonymous):

let me check my math again

OpenStudy (anonymous):

in any case i think this is increasing over your interval, and i don't see any .5 in the problem

OpenStudy (anonymous):

oh no i am wrong there is a critical point in there

OpenStudy (anonymous):

it is at \(t=\frac{361}{729}\) which is not really .5 but close enough i guess

OpenStudy (anonymous):

since you are working with annoying decimals anyway, might as well round

OpenStudy (anonymous):

how'd you get .81 instead of -.19? for your derivative

OpenStudy (anonymous):

you wrote \(.81t\) and the derivative of \(.81t\) is \(.81\)

OpenStudy (anonymous):

i was going to ask you where the \(-.19\) came from

OpenStudy (anonymous):

it was on the ans page

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

yw